OpenStudy (anonymous):

Which is the equation in slope-intercept form for the line that passes through (6, 16) and is perpendicular to 2x + 5y = 4?

OpenStudy (anonymous):

If the line is perpendicular to 2x + 5y = 4 then its slope is the negative reciprocal of the line whose equation is given.

OpenStudy (anonymous):

2x + 5y = 4 5y = 4 - 2x y = 4/5 - 2x/5 y = -2x/5 + 4/5 in slope-intercept form. The slope of this line is -2/5, and the negative reciprocal is 5/2, which would be the slope of any line perpendicular to this.

OpenStudy (anonymous):

Using this slope and the point given, you can find the equation of the second line.